# Learning Policies for Dynamic Coalition Formation in Multi-Robot Task Allocation

### v2

@@ V-B Simulation Results
Figure 4: The performance of the proposed framework is analyzed with respect to the task revision rate, i.e., the average number time steps that each robot changes its task selection, with data points collected over the last 500,000 training steps for each task spawn scenario.
@@
Models I, II, and III in this study are closely related to existing works [35], [17], and [23], respectively, but include the following necessary modifications: (i) All three models are trained using MAPPO (CTDE) within a Dec-POMDP framework, whereas the original references rely on the MDP assumption with individual rewards. (ii) While [17] focuses on individual-robot training scenarios, model II is trained for MRTA using our framework. (iii) In [23], only level-1 tasks are considered, and a method similar to Independent Q-Learning (IQL) [42], which encourages competition rather than cooperation among multiple robots, is used, while model III uses MAPPO.
@@
While frequent adjustments to task execution plans might seem disruptive to coalition formation among multiple robots working on high-level tasks, our proposed method enables robots to learn an optimal task revision strategy, ultimately resulting in superior performance. Figure 4 illustrates how our model effectively learns this strategy across various task spawn configurations. As shown, the performance of our framework tends to decline with excessive task assignment revisions. Conversely, without task revision, – corresponding to model III in the ablation study – the performance is significantly suboptimal, as illustrated in Fig. 3.
@@
In the case of a homogeneous task distribution, model I performs comparably to our model across all Bernoulli spawn settings but underperforms in the instant respawn scenario. This outcome indicates that homogeneous task scenarios with Bernoulli spawn do not necessitate long-term planning. However, in the instant respawn scenario, the constant task density discourages simplistic approaches like those employed by model I, resulting in its reduced performance.
@@ V-C Scalability and Generalizability Evaluation
Fig. 5(b) shows the generalizability scores of policies trained in easy, medium, and hard task settings, as well as in a combined setting encompassing all three. The results demonstrate that the policy trained across all settings achieves better and more consistent generalization. Notably, this policy performs close to the highest performance attainable by setting-specific policies. These results indicate that our method enables trained policies to allow robots to achieve near-optimal performance across environments with varying task-level distributions when trained in diverse environment settings.
@@ VI Conclusions
We introduced a learning-based framework to address decentralized dynamic coalition formation. By leveraging the MAPPO algorithm and incorporating spatial action maps, motion planning, task allocation revision, and intention sharing, our model demonstrated significant performance improvements over existing approaches. Extensive simulation studies highlighted the scalability of our framework with an increasing number of robots and its generalizability, performing well across diverse environments.
Future directions include improving the learning algorithm by integrating learned communication strategies in place of predefined intention-based communication. Additionally, we aim to explore mechanisms for achieving consensus among robots to expedite coalition formation, particularly for bandwidth-abundant environments.
Modifications to PCFA
We implemented PCFA and integrated it with GyMRTA for the performance comparison. To ensure a fair comparison with our method, the following modifications are applied: (i) Utility Function: The utility function $s_{k}$, associated with the $k$-th task, was adjusted to align with the reward function used in our framework: $s_{k}(t_{\text{ETA}})=l_{k}^{2}\,e^{-2(t_{\text{ETA}}+t_{w}-t_{k}^{0})},$ where $l_{k}$ is the task level, $t_{\text{ETA}}$ is the estimated time of arrival, $t_{w}$ is the wait time, and $t_{k}^{0}$ is the task’s initial spawn time. (ii) Partial Observability: In PCFA, the utility function for a task is set to zero if the task lies outside the robot’s view range $R_{\text{view}}$, preventing robots from forming coalitions for tasks they cannot observe. Additionally, the communication graph is updated at each time step, with links existing only between robots that are within the communication range $R_{\text{comm}}$. (iii) Task Assignment Revision: To align with our method, robots are allowed to revise their task assignments at every time step. This is achieved by recomputing all four stages of PCFA at every time step. While this modification increases communication overhead compared to the original PCFA, it allows a fair comparison in our simulation study.

### v3

@@ V-B Simulation Results
Figure 4: The graphs illustrate how task revision rates – defined as the average number of time steps between each robot’s task selection changes – evolve over the course of training, and how the average episodic reward improves as the robots learn optimal revision strategies. The dotted horizontal lines indicate the optimal revision rates.
@@
Models I, II, and III in this study are closely related to existing works [10], [11], and [12], respectively, but include the following necessary modifications: (i) All three models are trained using MAPPO within a Dec-POMDP framework, whereas the original references rely on the MDP assumption with individual rewards. (ii) While [11] focuses on individual-robot training scenarios, model II is trained for MRTA using our framework. (iii) In [12], only level-1 tasks are considered, and a method similar to Independent Q-Learning (IQL) [40], which encourages competition rather than cooperation among multiple robots, is used, while model III uses MAPPO.
@@
While frequent adjustments to task execution plans might seem disruptive to coalition formation among multiple robots working on high-level tasks, our proposed method enables robots to learn an optimal task revision strategy, resulting in superior performance. Fig. 4 depicts how the robots improve their average episodic reward by learning optimal task revision strategies. Importantly, this behavior is not manually encoded – specifically, the task revision rate is not a learnable parameter – but rather emerges from the learning process within our framework. Through repeated training, the robots progressively refine their optimal task revision rates, leading to more effective coalition formation and consistent improvements in average reward.
@@
In the homogeneous task distribution case, model I performs comparably to our model across all Bernoulli spawn settings but underperforms in the instant respawn scenario. This outcome indicates that homogeneous task scenarios with Bernoulli spawn do not necessitate long-term planning. However, in the instant respawn scenario, the constant task density discourages simplistic approaches like those employed by model I, resulting in its reduced performance.
@@ V-C Scalability and Generalizability Evaluation
Fig. 5(b) shows the generalizability scores of policies trained in easy, medium, and hard task settings, as well as in a combined setting encompassing all three. The results demonstrate that the policy trained across all settings achieves better and more consistent generalization. This policy performs close to the highest performance attainable by setting-specific policies. These results indicate that our method enables trained policies to allow robots to achieve near-optimal performance across environments with varying task-level distributions when trained in diverse environment settings.
@@ VI Conclusions
We introduced a learning-based framework to address decentralized dynamic coalition formation. By leveraging the MAPPO algorithm and incorporating spatial action maps, motion planning, task allocation revision, and intention sharing, our model demonstrated significant performance improvements over existing approaches. Future directions include conducting experiments with physical multi-robot systems, such as multiple mobile manipulators collaboratively transporting heavy objects, where sensor noise and communication delays may substantially impact the performance of the proposed framework. Additionally, we plan to explore fine-tuning our learning-based framework using data collected from these real-world experiments to further optimize its performance.
Modifications to PCFA: We implemented PCFA and integrated it with GyMRTA for the performance comparison. To ensure a fair comparison with our method, the following modifications are applied: (i) Utility Function: The utility function $s_{k}$, associated with the $k$-th task, was adjusted to align with the reward function used in our framework: $s_{k}(t_{\text{ETA}})=l_{k}^{2}\,e^{-2(t_{\text{ETA}}+t_{w}-t_{k}^{0})},$ where $l_{k}$ is the task level, $t_{\text{ETA}}$ is the estimated time of arrival, $t_{w}$ is the wait time, and $t_{k}^{0}$ is the task’s initial spawn time. (ii) Partial Observability: In PCFA, the utility function for a task is set to zero if the task lies outside the robot’s view range $R_{\text{view}}$, preventing robots from forming coalitions for tasks they cannot observe. Additionally, the communication graph is updated at each time step, with links existing only between robots that are within the communication range $R_{\text{comm}}$. (iii) Task Assignment Revision: To align with our method, robots are allowed to revise their task assignments at every time step. This is achieved by recomputing all four stages of PCFA at every time step. While this modification increases communication overhead compared to the original PCFA, it allows a fair comparison in our simulation study.